OpenStudy (anonymous):

Choose one of the factors of 5x3 – 135 5 x – 3 x2 + 3x + 9 All of the above

OpenStudy (jiteshmeghwal9):

take 5 common from the equation :)

OpenStudy (anonymous):

ooops make that \(5(x^3-27)\) as a start, then use \[a^3-b^2=(a-b)(a^2+ab+b^2)\] with \[a=x,b=3\]

OpenStudy (anonymous):

damn another typo, make that \[a^3-b^3=(a-b)(a^2+ab+b^2)\] with \(a=x\) and \(b=3\) just make a direct substitution into the formula for factoring the "difference of two cubes"

OpenStudy (anonymous):

@raebaby420 is it clear how to do that?